OpenStudy (anonymous):

solve for x: x/x+12 = 1/x-1

OpenStudy (anonymous):

Multiply both sides with (x+12)(x-1) so \[\Large (x)(x-1)=(x+12)\] \[\Large x^2-x=x+12\] \[\Large x^2-2x-12=0\] use Quadratic formula to solve this !

OpenStudy (anonymous):

OR you can try completing square !

hero (hero):

Whatever happened just simply cross multiplying?

OpenStudy (anonymous):

whats ur point? @Hero

hero (hero):

I already made my point @sami-21